OpenStudy (anonymous):

Given tan B = .34 find angle B in radians. Round your answer to the nearest hundredth.

OpenStudy (zale101):

\[B=\tan^{-1} (0.34)=?\]

OpenStudy (zale101):

Can you solve it from up here? Does this look familiar to you?

OpenStudy (anonymous):

Tangent Function:\[\large{\tan \theta=x~~~\implies~~~\theta=\tan^{-1}(x)}\]Degrees --> Radians:\[\large{\theta\times{\pi\over180}}\]